# A Survey of AI Methods for Geometry Preparation and Mesh Generation in Engineering Simulation

### v1

@@ 9 Parallel Mesh Generation
Positioned within this landscape, the surrogate-based study of [chrisochoides2025] addresses the parameter-tuning bottleneck by predicting execution time and enabling large synthetic searches that would otherwise consume weeks of HPC time. In parallel, other AI efforts intervene at structural decision points in the pipeline through learned load balancing [Gangopadhyay2025ParCFD], learned partitioning [Gatti2022RLGNN], and learning-based AMR [Yang2023RLAMR]. A near-term priority is portability for surrogate models beyond the training environment across geometry families and hardware by enriching inputs with quantitative geometric descriptors [foteinos2014, tsolakis2021] and by using transfer learning to adapt with minimal new data. A second priority is composition: pair surrogates with learned controllers to form an end-to-end AI-assisted workflow that partitions the domain, balances work, refines adaptively, and tunes runtime parameters under a single optimization loop. If realized, these advances would shift parallel meshing from a hand-tuned craft to a data-driven and reusable service embedded in large-scale simulation pipelines.
@@ 10 Scripting Automation
RL provides a natural way to address this challenge by framing script generation as a sequential decision-making problem. Unlike supervised methods that provide static recommendations, RL is designed to learn strategies that unfold over multiple steps. By treating CAD software as the environment, geometry operations as actions, and meshing or analysis outcomes as reward signals, RL agents can learn not only which operations to apply but also the order in which to apply them. Training on existing analyst scripts provides a foundation, while reward functions can encode expert priorities such as meshability, robustness, and fidelity to the target physics. In this way, RL has the potential to automate the generation of scripts that prepare CAD models for simulation, reducing manual effort and improving both efficiency and reproducibility.
Reinforcement learning methods for geometry preparation are implemented by casting CAD operations into the standard RL framework of environments, actions, and rewards. The environment is typically provided through an API to an existing geometry kernel, such as Cubit®  [CubitWebsite] or Open Cascade [OpenCascadeWebsite], which exposes the full set of modeling operations to the RL agent. Actions correspond to individual geometry commands that can be applied in sequence to modify the model, while the reward function encodes the analyst’s objectives, such as improving meshability or reducing unnecessary complexity. Training proceeds through trial-and-error interaction, allowing the agent to explore command sequences and gradually refine strategies that lead to high-quality model preparation scripts.
@@
Recent work has demonstrated the viability of this approach in several application areas. In the context of block decomposition for hexahedral meshing, DiPrete et al. [diprete2024reinforcement] and Zhang et al. [zhang2025reinforcement] designed RL agents whose actions correspond to decomposition operations. The agents learn to select reference entities and cutting operations that subdivide complex geometries into meshable sub-blocks. The reward functions in these studies were crafted to encourage the formation of well-shaped sub-blocks with favorable surface and dihedral angles, while penalizing irregular cuts that degrade mesh quality. To support effective training, Zhang et al. [zhang2025reinforcement] also introduced a graph-based representation of CAD models. This representation maps B-rep entities to NURBS-based graph nodes, enabling the neural network to capture geometric and topological relationships. Using this setup, they reported hex-meshable volume ratios exceeding 95%, a significant milestone for automated decomposition.
@@
Another example is the reduction of thin volumes in CAD models, a task that often requires ordered command sequences to collapse narrow solids into topologically valid sheet bodies. Owen et al. [owen2024cad] applied RL to this problem by defining actions that select thin volumes and determine appropriate sheet body placements. This approach preserved model topology while simplifying the geometry, and it is now being extended to automate defeaturing operations that improve tetrahedral mesh quality and reduce element counts. The long-term goal of this line of work is to provide analysts with automated script generation that mirrors expert decision-making, thereby reducing manual effort and accelerating the overall modeling-to-simulation workflow.
@@
Although these early demonstrations show considerable promise, the maturity of RL methods for geometry preparation is still at an early stage. Most studies remain proofs of concept, focusing on constrained problem classes such as block decomposition or thin volume reduction. These results confirm that RL can, in principle, learn sequences of CAD operations that prepare models for simulation, but the methods have not yet been validated across the wide range of geometries encountered in real engineering practice. One notable exception is the thin volume reduction work of Owen et al. [owen2024cad], which has been integrated into Cubit® and is already providing useful capability. Even here, however, broader training data and wider analyst usage are needed before the approach can be considered mature.
@@
Several strategies are being explored to address this issue. One approach is to convert B-rep entities into structured formats that can be processed by neural networks, such as sequence-based encodings [zhang2024brep2seq, zhang2024brepmfr]. Another is to use graph-based representations, where nodes and edges capture both geometry and topology, and graph neural networks (GNNs) are applied to process this information [wu2020comprehensive]. Graph-based methods are particularly promising because they allow the RL agent to reason about how operations modify geometry and topology in ways directly relevant to meshing and analysis. By combining these representation strategies with reinforcement learning, it may be possible to train agents that not only reproduce expert scripts but also discover new and more efficient procedures for geometry preparation.
@@ 11 Language-Driven Scripting
While reinforcement learning offers a framework for discovering sequences of CAD operations through trial and error interaction with geometry software, large language models (LLMs) [zhao2023survey] address a complementary aspect of the problem: the automation of CAD scripting through direct code synthesis. RL methods are designed to learn strategies over many iterations, discovering which procedures lead to successful outcomes. In contrast, LLMs are trained on large collections of text and code, enabling them to directly translate natural language prompts into executable scripts. Within the broader context of geometry preparation and meshing, these two approaches are not competing but rather mutually reinforcing. RL develops new procedural strategies, while LLMs provide an accessible interface for capturing, reusing, and deploying these strategies in a reproducible manner.
@@
The motivation for LLM-driven scripting stems from the same challenges that have long motivated automation: CAD workflows require careful, reproducible documentation of geometry operations. Analysts traditionally create scripts manually, either through tedious trial and error in a graphical interface or by directly writing commands. While effective, this practice is slow, error-prone, and demands specialized expertise. LLMs reduce these barriers by allowing analysts to describe their goals in natural language. From these descriptions, the model generates CAD commands that can reproduce geometry, apply defeaturing, initiate meshing procedures, or prepare models for analysis. Importantly, the user remains in control: the generated code is transparent and can be inspected, refined, or overridden as needed.
@@
Recent developments highlight several themes in how LLMs are being applied to CAD scripting. One line of work focuses on direct code generation. Fine-tuned models and database retrieval-augmentation frameworks have shown the ability to translate textual input into syntactically correct and semantically meaningful CAD code [R1, R4, R13]. These approaches often employ structured function calling or specialized annotation schemes that capture the semantics of CAD operations such as extrusions, fillets, and lofts [N1]. Iterative refinement loops, whether automated or assisted by user feedback, further improve robustness by detecting errors and resynthesizing code [R5, R6, R14].
@@
A second direction expands beyond text to multimodal input. Systems now combine language with sketches, images, or point clouds, enabling workflows where a designer provides partial visual guidance that the model translates into complete parametric scripts. Examples include models trained on large datasets of CAD images paired with code [R13, R18, R26], as well as fine-tuned multimodal LLMs that directly edit boundary representations from text prompts while preserving geometric integrity [N2]. These methods move toward richer interactions where LLMs act as design assistants capable of interpreting intent expressed across multiple modalities.
A third theme emphasizes verification and correction. Because CAD scripting demands high fidelity, LLM-based workflows frequently incorporate feedback mechanisms. Some systems compare generated geometry against target specifications, iteratively correcting deviations [R5]. Others employ auxiliary vision–language models to check code execution and flag errors for resynthesis. These feedback loops echo the trial and error philosophy of RL, but they are carried out in a supervised or user-guided manner that accelerates convergence toward reliable scripts.
@@
Together, these advances show that LLMs complement the procedural learning strengths of RL by providing a unifying interface that connects geometry representations, meshing technologies, and AI-driven methods discussed throughout this paper. Where RL agents may learn how to optimally defeature a model or decompose it for hexahedral meshing, an LLM can package such procedures into user-accessible scripts. Representations such as graph neural networks or B-rep encodings, originally motivated by the need to support RL training [zhang2024brep2seq, zhang2024brepmfr, wu2020comprehensive], can also support LLMs by enriching the context available for script generation.
@@
Looking forward, the convergence of these technologies points toward the rise of AI agents [fang2025selfevolving] for CAD and meshing. Agents extend the capabilities of LLMs by combining language understanding with iterative reasoning, tool usage, and memory. In the context of geometry preparation, such agents could autonomously call geometry kernels, verify outputs, and refine operations, while engaging users through natural language. This unifies the strengths of both paradigms: RL provides the procedural foundation by discovering effective sequences of operations, LLMs make those procedures accessible by generating scripts from high-level prompts, and agent architectures orchestrate the entire process in a closed loop. The result is a vision of intelligent assistants that accelerate analysis, ensure reproducibility, and lower barriers to entry, while keeping analysts firmly in control of the workflow.
Recent advances indicate that LLM-driven CAD scripting systems perform reliably for routine design tasks [R9, R10, R13, R14]. Nonetheless, challenges persist. Models may struggle with highly constrained or complex geometries, and iterative refinement remains necessary to address error propagation and maintain high fidelity in the designs [R15, R17, R28, R30]. These limitations lead the literature to highlight areas that should be prioritizes as future research:
@@
  • Expanding Domain-Specific and Multimodal Datasets: The lack of large-scale, high-quality CAD-specific and multimodal datasets (text, sketches, images, and CAD programs) limits the ability of LLMs to generalize across diverse design tasks. Future research should focus on curating and releasing broader datasets that capture complex geometries and real-world industrial contexts [R2, R9, R18].
@@
  • Improving Spatial Reasoning and Geometric Fidelity: While LLMs can produce syntactically valid CAD scripts, they often fail on tasks requiring precise geometric or spatial reasoning. Developing solver-aided approaches, neurosymbolic frameworks, or enhanced spatial localization mechanisms could significantly improve fidelity for complex and constrained designs [R10, R23, R25].
@@
  • Extending LLMs Beyond Geometry to Full Design-to-Manufacturing Pipelines: Research should go beyond geometry generation to encompass other stages of engineering workflows, including design evaluation, optimization, feature tagging for simulation/FEA, and integration with downstream manufacturing processes. This would enable seamless end-to-end AI-assisted design systems [R12, R19, R20].
@@
In addressing these limitations, researchers can accelerate the evolution of LLMs to achieve a transformative shift in the digital design cycle. By automating the translation of natural language into precise and executable CAD code, these systems promise to accelerate the design process while lower the threshold for broad adoption.
@@ 12 Conclusion
Across these topics, several cross-cutting themes emerge. Most approaches face common bottlenecks: the lack of large curated datasets, standardized benchmarks, and robust geometry representations that can transfer across diverse CAD sources. The maturity of methods also varies widely. Classical machine learning for classification and mesh quality prediction has already entered production software such as Cubit®, delivering measurable efficiency gains. In contrast, reinforcement learning for meshing and volumetric parameterization remain at the prototype stage, while segmentation and defeaturing tools are intermediate—promising in research settings but not yet broadly deployed. The most promising direction is integration: connecting classification, prediction, and defeaturing into closed-loop workflows; combining supervised and reinforcement learning; and leveraging large language models for scripting and natural interfaces. Progress in the next five years will depend less on isolated algorithmic advances than on building shared datasets, evaluation standards, and hybrid pipelines that unify AI with established meshing practice.
@@
Acknowledgments Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525.

### v2

@@ 10 Parallel Mesh Generation
Together, these results suggest a practical division of labor: surrogates accelerate search over interdependent runtime parameters, while learned controllers directly optimize decomposition, load balancing, and adaptive refinement. A near-term priority is portability for surrogate models beyond the training environment across geometry families and hardware. This will be achieved by enriching the input with quantitative geometric descriptors [Foteinos_JPDC_2014, tsolakis2021, tsolakis2024] and using transfer learning to adapt with minimal new data. A second priority is composition, pairing surrogates with learned controllers to form an end-to-end AI-assisted workflow that partitions the domain, balances work, refines adaptively, and tunes runtime parameters under a single optimization loop. If realized, these advances would shift parallel meshing from a hand-tuned craft to a data-driven and reusable service embedded in large-scale simulation pipelines.
N. Chrisochoides acknowledges support from The Richard T. Cheng Endowment and thanks Kevin Garner and Min Dong for assistance with data curation and statistical analysis. Generative AI tools (Gemini, Grammarly) were used for language editing; the author reviewed and assumes full responsibility for the final content.
Contributed by N. Chrisochoides
@@ 11 Scripting Automation
Reinforcement learning (RL) provides a natural way to address this challenge by framing script generation as a sequential decision-making problem. By treating CAD software as the environment, geometry operations as actions, and meshing or analysis outcomes as reward signals, RL agents can learn not only which operations to apply but also the order in which to apply them. In practice, the environment is typically provided through an API to an existing geometry kernel, such as Cubit [CubitUserDoc1706_2025] or Open Cascade [OpenCascadeWebsite], which exposes modeling operations to the agent. Training proceeds through trial-and-error interaction and can be initialized from existing analyst scripts, while reward functions encode priorities such as meshability, robustness, and fidelity to the target physics. In this way, RL has the potential to automate the generation of scripts that prepare CAD models for simulation, reducing manual effort and improving both efficiency and reproducibility.
@@
Recent work has demonstrated the viability of this approach in several application areas. In the context of block decomposition for hexahedral meshing, DiPrete et al. [diprete2024reinforcement] and Zhang et al. [zhang2025reinforcement] designed RL agents whose actions correspond to decomposition operations. The agents learn to select reference entities and cutting operations that subdivide complex geometries into meshable sub-blocks. The reward functions in these studies were crafted to encourage the formation of well-shaped sub-blocks with favorable surface and dihedral angles, while penalizing irregular cuts that degrade mesh quality. To support effective training, Zhang et al. [zhang2025reinforcement] also introduced a graph-based representation of CAD models. This representation maps B-rep entities into graph structures that capture geometry and topology, enabling the neural network to reason about relationships that matter for decomposition. Using this setup, they reported hex-meshable volume ratios exceeding 95% on their benchmark problems, a notable milestone for automated decomposition.
@@
Another example is the reduction of thin volumes in CAD models, a task that often requires ordered command sequences to collapse narrow solids into topologically valid sheet bodies. Owen et al. [owen2024cad] applied RL to this problem by defining actions that select thin volumes and determine appropriate sheet body placements. Their framework learns a policy to drive these ordered operations within a CAD kernel, using per-volume agents and a connectivity graph to coordinate decisions across the model. This approach preserves model topology while simplifying the geometry, and it is now being extended to automate defeaturing operations that improve tetrahedral mesh quality and reduce element counts. The long-term goal of this line of work is to provide analysts with automated script generation that mirrors expert decision-making, thereby reducing manual effort and accelerating the overall modeling-to-simulation workflow.
A key practical question is how to evaluate generated scripts in ways that reflect engineering needs. Beyond task completion, relevant criteria include meshing success rates, mesh quality distributions, element count and resolution relative to target physics, robustness under small geometric variations, and the degree to which scripts preserve functional and interface features. Related to this, real workflows impose hard constraints, for example preserving watertightness, respecting tolerances, and avoiding topology changes that invalidate boundary conditions. Incorporating these constraints into the action space and reward design is often necessary for RL policies to be deployable in production settings.
@@
Although these early demonstrations show considerable promise, the maturity of RL methods for geometry preparation is still at an early stage. Most studies remain proofs of concept, focusing on constrained problem classes such as block decomposition or thin volume reduction. These results confirm that RL can, in principle, learn sequences of CAD operations that prepare models for simulation, but the methods have not yet been validated across the wide range of geometries encountered in real engineering practice. One notable exception is the thin volume reduction work of Owen et al. [owen2024cad], which has been integrated into Cubit [CubitUserDoc1706_2025] and is already providing useful capability. Even here, broader training data, clearer success criteria, and wider analyst usage are needed before the approach can be considered mature. More broadly, generalization remains challenging due to CAD heterogeneity, kernel and version differences, tolerance-sensitive predicates, and long-horizon workflows where rewards are sparse.
@@
Several strategies are being explored to address this issue. One approach is to convert B-rep entities into structured formats that can be processed by neural networks, such as sequence-based encodings [zhang2024brep2seq, zhang2024brepmfr]. Another is to use graph-based representations, where nodes and edges capture both geometry and topology, and graph neural networks (GNNs) are applied to process this information [wu2020comprehensive]. Graph-based methods are particularly promising because they allow the RL agent to reason about how operations modify geometry and topology in ways directly relevant to meshing and analysis. By combining these representation strategies with reinforcement learning, it may be possible to train agents that not only reproduce expert scripts but also discover new and more efficient procedures for geometry preparation. Looking ahead, these RL capabilities are also complementary to language-model-based scripting, where natural-language intent can propose candidate procedures and RL can refine them through closed-loop interaction with the geometry kernel.
Contributed by N. Winovich
@@ 12 Language-Driven Scripting
While reinforcement learning offers a framework for discovering sequences of CAD operations through trial and error interaction with geometry software, large language models (LLMs) [zhao2023survey] address a complementary aspect of the problem: the automation of CAD scripting through direct code synthesis. RL learns strategies over many iterations; LLMs translate natural language prompts into executable scripts. Within geometry preparation and meshing, these approaches are mutually reinforcing. RL develops procedural strategies, while LLMs provide an interface for capturing, reusing, and deploying these strategies in reproducible form.
@@
The motivation for LLM-driven scripting stems from the same challenges that have long motivated automation: CAD workflows require careful, reproducible documentation of geometry operations. Analysts traditionally create scripts manually, either through tedious trial and error in a graphical interface or by directly writing commands. While effective, this practice is slow, error prone, and demands specialized expertise. LLMs lower these barriers by allowing analysts to describe goals in natural language. From these descriptions, the model generates CAD commands that reproduce geometry, apply defeaturing, initiate meshing procedures, or prepare models for analysis. The user remains in control because the generated code is transparent and can be inspected, refined, or overridden as needed.
@@
Three themes characterize current applications of LLMs to CAD scripting. First, direct code generation uses fine-tuned models and database retrieval-augmentation frameworks to translate text into syntactically correct and semantically meaningful CAD code [R1, R4, R13]. These approaches often employ structured function calling or specialized annotations that capture the semantics of operations such as extrusions, fillets, and lofts [N1]. Iterative refinement, whether automated or guided by user feedback, improves robustness by detecting errors and resynthesizing code [R5, R6, R14].
@@
Second, multimodal input extends beyond text by combining language with sketches, images, or point clouds, enabling workflows in which a designer supplies partial visual guidance that the model converts into complete parametric scripts. Examples include models trained on large datasets of CAD images paired with code [R13, R18, R26], as well as fine-tuned multimodal LLMs that directly edit boundary representations from text prompts while preserving geometric integrity [N2]. These methods move toward richer interactions where LLMs act as design assistants that interpret intent across sketches, images, and point clouds and translate it into complete parametric scripts.
Third, verification and correction are integral because CAD scripting demands high fidelity. Many workflows compare generated geometry against target specifications and iteratively correct deviations [R5]. Others use auxiliary vision–language models to check code execution and flag errors for resynthesis. Menon et al. present a multi-agent design assistant (MADA) [Menon2026MultiAgentISC] that exemplifies this theme by grounding command synthesis in documentation retrieval, checking mesh metrics, and looping through scheduler-managed runs to converge on target quantities of interest.
In MADA, the agentic workflow is organized around complementary roles. A Geometry Agent creates and edits CAD, invokes defeaturing utilities, and generates meshes. A Job Agent manages scheduler interaction to launch and monitor ensembles at scale. An Inverse Design Agent explores parameters using simulation outputs or surrogate models. This division mirrors how engineering teams split modeling, meshing, and analysis, and it illustrates how LLM-driven scripting and RL-style search can be orchestrated within a single closed loop [Menon2026MultiAgentISC]. Related work extends language-driven workflows beyond CAD code to mesh artifacts themselves. LLaMA-Mesh [wang2024llamamesh] encodes mesh geometry and connectivity in a text-compatible form, enabling instruction-driven mesh generation and editing that can support similar verification and refinement loops in mesh-centered pipelines.
@@
Together, these advances show that LLMs complement the procedural learning strengths of RL by providing a unifying interface that connects geometry representations, meshing technologies, and AI-driven methods discussed throughout this paper. Where RL agents may learn how to optimally defeature a model or decompose it for hexahedral meshing, an LLM can package such procedures into user-accessible scripts. Representations such as graph neural networks or B-rep encodings, originally motivated by the need to support RL training [zhang2024brep2seq, zhang2024brepmfr, wu2020comprehensive], can also support LLMs by enriching the context used during script generation.
@@
Looking forward, the convergence of these technologies points toward AI agents [fang2025selfevolving] for CAD and meshing. Agents combine language understanding with iterative reasoning, tool usage, and memory. In geometry preparation, such agents could autonomously call geometry kernels, verify outputs, and refine operations while engaging users through natural language. This unifies the strengths of both paradigms: RL provides procedural building blocks, LLMs make those procedures accessible through high-level prompts, and agent architectures orchestrate the process in a closed loop. Experience with the assistant suggests that the pattern scales to realistic CAD-to-mesh-to-solver use cases and to HPC settings [Menon2026MultiAgentISC].
@@
Expanding domain-specific and multimodal datasets: The lack of large-scale, high-quality CAD-specific and multimodal datasets limits generalization. Priority should be placed on broader datasets that capture complex geometries and industrial contexts [R2, R9, R18].
@@
Improving spatial reasoning and geometric fidelity: While LLMs can produce syntactically valid CAD scripts, precision remains challenging for tightly constrained designs. Solver-aided methods, neurosymbolic frameworks, or stronger spatial localization could improve fidelity [R10, R23, R25].
@@
Extending beyond geometry to full design-to-manufacturing pipelines: Research should incorporate design evaluation, optimization, feature tagging for simulation and FEA, and downstream manufacturing integration to enable end-to-end AI-assisted design systems [R12, R19, R20].
@@
In addressing these limitations, researchers can accelerate the evolution of LLMs to achieve a transformative shift in the digital design cycle. By automating the translation of natural language into precise and executable CAD code, these systems promise to accelerate the design process while lowering the threshold for broad adoption.
Contributed by N. Brown
@@ 13 Conclusion
Across these topics, several cross-cutting themes emerge. Most approaches face common bottlenecks: the lack of large curated datasets, standardized benchmarks, and robust geometry representations that can transfer across diverse CAD sources. The maturity of methods also varies widely. Classical machine learning for classification and mesh quality prediction has already entered production software such as Cubit [CubitUserDoc1706_2025], delivering measurable efficiency gains. In contrast, reinforcement learning for meshing and volumetric parameterization remain at the prototype stage, while segmentation and defeaturing tools are intermediate—promising in research settings but not yet broadly deployed. The most promising direction is integration: connecting classification, prediction, and defeaturing into closed-loop workflows; combining supervised and reinforcement learning; and leveraging large language models for scripting and natural interfaces. Progress in the next five years will depend less on isolated algorithmic advances than on building shared datasets, evaluation standards, and hybrid pipelines that unify AI with established meshing practice.
@@
Acknowledgments
Chrisochoides is funded by The Richard T. Cheng Endowment. He thanks Dr. Kevin Garner and Min Dong assisted in curating the data for both shared memory (NUMA) and distributed memory parallel mesh generation runs at CRTC and statistics from the wait-time of those runs at ODU’s HPC Clusters, respectively.
The authors used ChatGPT, Gemini and Gramerly for language editing and readability improvements. All authors reviewed the manuscript and take full responsibility for the final content.